Laser Absorption Spectroscopy of LaF: Analysis of the B^{1}Π- X^{1}Σ^{+} Transition

L. A. Kaledin,\underline{A. L. Kaledin}, and M. C. Heaven, Department of Chemistry, Emory University, Atlanta, GA 30322.

Laser absorption spectrum of the LaF B^{1}Π-X^{1}Σ^{+} (2,0), (1,0), (2,1), and (3,2) bands was recorded. Gas phase LaF was obtained by heating a mixture of a 0.2 g of La metal and 0.3 g of LaF_{3} powder to 1900 K. In order to modulate the absorption signal, the concentration of LaF was varied by passing a 25 kHz discharge (80-100W, 0.5A peak current) through the vapor. Under these conditions the Tantalum heating tube used to vaporize the sample also acted as a hollow cathode. The B-X band system was observed previously ^{a}. In analyzing the spectrum, we have combined absorption data with that from the Reference. Accurate molecular constants (in cm^{−1}) and internuclear distance r_{e} (in nm) for the X^{1}Σ^{+} state are:

ω_{e} = 573.318, ω_{e}x_{e} = 2.18, B_{e} = 0.246411(15), α_{e} = 1.251(15) × 10^{−3}, r_{e} = 0.20232.

Rotational perturbations were evident in the B^{1}Π v′ =1,v′ =2,and v′ =3 levels. The pattern of the B^{1}Π state perturbations was shown to be a consequence of the interaction with the c_{1}^{3}Δ_{1} state by using the experimental value of ω_{e} =450 cm^{−1} (see Reference) and estimated value of ω_{e}x_{e} =3.5 cm^{−1} for the triplet state. The principal new results to emerge from this investigation are the determination of the equilibrium constants for the B^{1}Π and X^{1}Σ^{+} states of LaF.

Work supported by AFOSR under grant F19628-90-C-025.

aH. Schall, C. Linton, and R. W. Field, J. Mol. Spectrosc. 100, 437-448 (1985); B. Simard and A. M. James, J. Chem. Phys.

97,4669-4678 (1992)